OpenStudy (anonymous):

What is the solution to the rational equation 2x/3 - 1/3 = 5

OpenStudy (anonymous):



OpenStudy (anonymous):

Options are.... x = 8 x = 3 x = 7 x = 11

OpenStudy (anonymous):

2x -1 = 3 * 5 = 15

OpenStudy (anonymous):

Whats 4+4?